[Brahms-l] Shift report 20040214 08:00-16:00

From: Apache <apache@bnl.gov>
Date: Sat Feb 14 2004 - 15:59:36 EST
------------------------------------------------------------
Date: Sat Feb 14 15:59:36 2004, Shift: 08:00-16:00
Supervisor: DO, with: 
Log book#15 pages 57 - 57

Shift Summary:
------------------------------------------------------------
1.Nice stroes (a problem with a power supply control 
  in RHIC between 2 stores).

2.Data runs:     
           FS 4d B1/1    MRS 45d B1050 

               tr2           tr3

    10338      44k           63k
    10339      60k          126k 
    10342      56k           73k 
    10343      67k          100k 
    10344     132k          150k 
    10345      61k           85k
    10346      56k           81k              
    10347      still taking data
 
  Pedestals: non TPC 10340,  TPC 10341

3.BRAHVO cannot print out to qms_ch, I had to transfer log
  files from pii4 to another machine and print directly
  from there. Supermon can print again.
  Detectors behave very good. Note that TRFS top tdc 7 has
  resurrected for a bunch of runs but is dead again.
     
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sat Feb 14 15:59:56 2004

This archive was generated by hypermail 2.1.8 : Sat Feb 14 2004 - 16:00:28 EST